# Rethinking Cross-Attention for Infrared and Visible Image Fusion

### v1

@@ IV-C Results and Discussion
2) Results on MSRS Dataset: Fig. 7 displays the fusion results obtained by the seven comparison methods on the MSRS dataset. Although the existing fusion methods can obtain results by retaining information from the source images to some extent, they still encounter detail blurring and salient information loss. Specifically, the results generated by DATFuse lose different degrees of detail information from the visible image (see Fig. 7(c)). DenseFuse, FusionGAN, RFN-Nest, and U2Fusion fail to effectively balance the infrared information and visible information, such as the blurred effects around the tower edges in the red boxes of the last row. SwinFusion seems to have richer details and brightness information. However, it loses some details from the infrared images, since this method utilizes a pixel maximum constraint loss for each pixel. (see the red boxes in the first and fifth rows). In contrast, our ATFuse achieves a better tradeoff between detail preservation and salient information retention.
@@
TABLE I: Using six metrics to quantitatively compare ATFuse with six existing fusion methods on the MSRS dataset (twenty image pairs). The best and the second best performance is marked by bold and underlined, respectively.
Methods $AG$ $EN$ $SD$ $SF$ $VIFF$ $Qabf$
DATFuse 4.1944 6.8679 9.3450 0.0529 0.8020 0.5919
DenseFuse 2.5050 6.4331 8.2278 0.0284 0.7052 0.3733
FusionGAN 1.7466 5.7239 6.6471 0.0200 0.5730 0.1358
RFN-Nest 2.6218 6.6975 8.5525 0.0282 0.7499 0.4152
SwinFusion 4.3921 7.1548 9.1247 0.0526 1.0195 0.6332
U2Fusion 2.5714 6.3092 7.9221 0.0297 0.6607 0.3920
Ours 4.6872 7.1441 8.9113 0.0551 0.9718 0.6982
Table I displays objective evaluations of seven different fusion methods on twenty image pairs from the MSRS dataset. Our ATFuse method performs best in the $AG$, $SF$, and $Q_{abf}$ metrics. The remaining three metrics are slightly different from the best ones. Specifically, our method performs second-best in the $EN$ and $VIFF$ metrics and third-best in the $SD$ metric. Overall, based on subjective and objective analysis, we can conclude that our ATFuse method achieves the best fusion performance and outperforms other methods in both visual perception and objective metrics.
@@
3) Extension Experiment on TNO Dataset: To further validate the effectiveness and generalization capabilities of our method, we directly utilize the trained model to test the generalization on the TNO dataset. Fig. 8 shows five image pairs from the TNO dataset and the corresponding fusion results obtained by seven methods. We have magnified two local areas in each image for better comparison. It can be seen that the six comparison methods still have some drawbacks compared to our method. Specifically, DenseFuse, RFN-Nest, and U2Fusion fail to extract enough salient information from the infrared images compared to the other methods (see the red box in the second row and the green box in the fourth row). Moreover, the results produced by FusionGAN are mostly blurry around the edges (see the green box in the fourth row and the red box in the fifth row). For DATFuse, SwinFusion, and our method, the overall saliency in DATFuse is notably worse (see the green box in the fourth row), because it adopts an average constraint approach for each pixel in the loss function, leading to a saliency reduction of the results. SwinFusion loses some detailed information in the visible image due to its loss function making the values of all pixels in the fused image close to the maximum values of the corresponding pixels in the two source images (see the red box in the fifth row). In contrast, the proposed segmented pixel loss function makes the fused pixel values tend toward the maximum values of corresponding pixel points in target regions, resulting in better preservation of salient information. Furthermore, our method effectively extracts discrepancy information from two source images, whereas SwinFusion and DATFuse struggle to capture the distinct information of infrared images (see the green box in the third row).
TABLE II: Using six metrics to quantitatively compare ATFuse with six existing fusion methods on the TNO dataset (twenty image pairs). The best and the second best performance is marked by bold and underlined, respectively.
Methods $AG$ $EN$ $SD$ $SF$ $VIFF$ $Qabf$
DATFuse 3.1016 6.3137 8.1289 0.0334 0.7214 0.4802
DenseFuse 2.3471 6.1386 7.9529 0.0236 0.6065 0.3435
FusionGAN 2.2010 6.3553 8.0122 0.0226 0.6515 0.2190
RFN-Nest 2.7671 6.8298 8.7919 0.0243 0.7970 0.3649
SwinFusion 3.9429 6.7155 8.6421 0.0400 0.8467 0.5150
U2Fusion 2.4238 6.1319 7.9008 0.0226 0.6020 0.3593
Ours 4.3232 6.8774 8.5891 0.0431 0.8559 0.5900
@@
Table II reports that our method achieves the best fusion effects in $AG$, $EN$, $SF$, $VIFF$, and $Q_{abf}$ metrics on the TNO dataset. The $SD$ metric record of our method is third-best, with a small difference compared to the best performance. In general, the ATFuse method achieves the best objective and subjective performance among the seven methods, testing the twenty image pairs from the TNO dataset, and affirming its excellent generalization ability.
IV-D Ablation Studies
@@
In this section, we have conducted two ablation studies to investigate the effectiveness of our method, namely, network structures and the segmented pixel loss function.
@@
Figure 9: Fusion results obtained by three different network structures. (a) infrared image, (b) visible image, (c) the results of w/o DIIM, (d) the results of w/o ACIIM, (e) Ours.
@@
1) Ablation Study on Network Structure: To investigate the effectiveness of DIIM and ACIIM, we separately remove the DIIM (named “w/o DIIM”), and the ACIIM (named “w/o ACIIM”). We compare the fusion results obtained by w/o ACIIM, w/o DIIM, and the proposed method, as shown in Fig. 9. It can be seen that fusion results produced by using w/o DIIM only contain detail information from the visible image, while the salient information from the infrared image is almost entirely lost (see the green boxes in Fig. 9(c)). Besides, our fusion results have more detail information than the results by using w/o ACIIM network (see the red boxes Fig. 9(d-e)). The main reasons lie in that the proposed ATFuse embeds the DIIM and ACIIM modules, which have a good ability to capture the discrepancy information and the common information of the source images.
@@
TABLE III: Quantitative comparison of fusion results obtained by three different network structures. The best and the second best performance is marked by bold and underlined, respectively.
Methods $AG$ $EN$ $SD$ $SF$ $VIFF$ $Q_{abf}$
w/o DIIM 4.4747 6.9185 9.4503 0.0422 0.9755 0.4105
w/o ACIIM 5.3814 7.1582 10.2160 0.0545 0.8026 0.6123
Ours 5.6864 7.2355 10.2413 0.0572 0.8374 0.6295
@@
Table III provides an objective comparison of our ATFuse with another two network structures. Apparently, our method can achieve the best performance in $AG$, $EN$, $SD$, $SF$, $Q_{abf}$, and the second-best performance in $VIFF$. Therefore, the proposed DIIM and ACIIM can fully extract useful features, which enhance the retention of discrepancy information and the injection of common information in the fusion results, respectively.
@@
2) Ablation Study on Pixel Loss Function: To identify the impact of the segmented pixel loss function, we conduct four experiments by using different pixel loss functions $\mathcal{L}_{1}$ ($\alpha=100$), $\mathcal{L}_{2}$ ($\alpha=0$), $\mathcal{L}_{3}$ ($\alpha=50$) and $\mathcal{L}_{4}$ ($\alpha=80$) in Eq. (16). Actually, $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ correspond to pixel maximum constraint and average constraint, respectively. Fig. 10 shows the fusion outcomes of two image pairs with the five different loss functions. From the red boxes in Fig. 10 (c, e, f), it can be obviously seen that the results generated by using $\mathcal{L}_{1}$, $\mathcal{L}_{3}$ and $\mathcal{L}_{4}$ have different degrees of detail loss. Besides, as the value of $\alpha$ increases, the degree of detail loss becomes serious. It is because of the increasing value of $\alpha$, more pixels in the image are constrained by the maximum value selection loss, which leads to some necessary information loss from one of the source images in fusion results. The results obtained by the $\mathcal{L}_{2}$ loss lack brightness information than the results of other loss functions (see the green box in Fig. 10(d)), this is because the essence of $\mathcal{L}_{2}$ loss is pixel average constraint. On the whole, setting $\alpha$ to 20 results in a satisfactory trade-off between maintaining detail and preserving saliency.
@@
TABLE IV: Quantitative comparison of fusion results obtained by the proposed method with five different pixel loss functions. The best and the second best performance is marked by bold and underlined, respectively.
Loss Functions $AG$ $EN$ $SD$ $SF$ $VIFF$ $Q_{abf}$
$\mathcal{L}_{1}$ ($\alpha=100$) 5.4194 7.1378 10.1297 0.0546 0.8298 0.5913
$\mathcal{L}_{2}$ ($\alpha=0$) 5.7090 7.1636 9.9904 0.0633 0.7897 0.6233
$\mathcal{L}_{3}$ ($\alpha=50$) 5.5444 7.1843 10.2633 0.0556 0.8281 0.6038
$\mathcal{L}_{4}$ ($\alpha=80$) 5.4129 7.1334 10.1725 0.0549 0.8391 0.6071
Ours ($\alpha=20$) 5.6864 7.2355 10.2413 0.0572 0.8374 0.6295
@@
Table IV presents the objective comparison of our ATFuse approach integrated with various loss functions. The average outcomes of six objective evaluation metrics are tabulated in Table IV. The best and second-best results for each metric are bolded and underlined, respectively. It can be seen results generated by our method outperform those of other loss functions in terms of most objective evaluations. Therefore, the choice of $\alpha$ to 20 is reasonable in our study.
IV-E Analysis of Some Configurations
@@
Figure 11: Fusion results obtained by the proposed method with different $\gamma$ values. (a) infrared image, (b) visible image, (c-e) the qualitative results of different $\gamma$.
@@
TABLE V: Quantitative comparison of fusion results obtained by different $\gamma$ values. The best and the second best performance is marked by bold and underlined, respectively.
Hyperparameter $AG$ $EN$ $SD$ $SF$ $VIFF$ $Q_{abf}$
$\gamma=0.5$ 5.3458 7.1918 10.2105 0.0528 0.8091 0.5727
$\gamma=0.75$ 5.4344 7.1711 10.2065 0.0552 0.8286 0.6212
$\gamma=1.0$ 5.6864 7.2355 10.2413 0.0572 0.8374 0.6295
@@
1) Hyperparameter $\gamma$: As mentioned in Section III-C, the hyperparameter $\gamma$ controls the weight of pixel loss and texture loss in the loss function. Based on our experience, we set $\gamma$ to 0.5, 0.75, and 1.0. Fig. 11 displays the fusion images produced by ATFuse using different $\gamma$ values in the loss function. It can be seen that the output obtained with $\gamma=1.0$ exhibits clear texture details around the windows. Table V presents an objective comparison of ATFuse using different $\gamma$ values. It is evident from the Table that setting $\gamma$ to 1.0 yields the best objective performance.
@@
TABLE VI: Quantitative comparison of fusion results obtained by the proposed method with different number of feature fusion modules. The best and the second best performance is marked by bold and underlined, respectively.
Number $AG$ $EN$ $SD$ $SF$ $VIFF$ $Q_{abf}$
Three 5.6557 7.2357 10.2076 0.0570 0.8431 0.6229
Two 5.4366 7.1983 10.1812 0.0558 0.8538 0.6215
One (Ours) 5.6864 7.2355 10.2413 0.0572 0.8374 0.6295
@@
2) Number of Feature Fusion Modules: We assembled one DIIM and two ACIIMs into a single feature fusion module, as shown in Fig. 1 The number of feature fusion modules can be decided by inspecting the fusion performance through objective metrics. Table VI reports the quantitative fusion results by using various number of feature fusion modules. It can be seen that using one feature fusion module in the proposed framework achieves the best overall effect. Therefore, considering computational efficiency, one group of modules is suitable to fuse features.
IV-F Computational Efficiency
TABLE VII: Average running time of one image pair on four different datasets by using different fusion methods (Unit: second). The best and the second best performance is marked by bold and underlined, respectively.
Datasets DATFuse DenseFuse FusionGAN RFN-Nest SwinFusion U2Fusion Ours
TNO (s) 0.1354 0.1066 0.2382 0.1781 1.6576 0.0675 0.1026
RoadScence (s) 0.1248 0.1082 0.3129 0.1515 0.9637 0.0682 0.1081
MSRS (s) 0.1419 0.1033 0.1401 0.1832 1.7043 0.0691 0.1054

### v2

@@ 4.3 Results and Discussion
1) Results on the MSRS Dataset: Fig. 5 displays the fusion results on the MSRS dataset. Although the existing fusion methods can obtain results by retaining information from the source images to some extent, they still encounter detail blurring and salient information loss. Specifically, DenseFuse, SwinFuse, RFN-Nest, U2Fusion, and LRRNet fail to preserve well brightness information from the infrared image in the fusion results (see the red box in the second column and the green box in the fourth column). The results generated by AEFusion lose different degrees of detail information, causing blurriness in some areas, see the red box in the third column and the green box in the fifth column. The results of DATFuse also exhibit poor performance in terms of texture information injection of the visible image and thermal target preservation of the infrared image, see the green box in the third column and the red box in the fourth column. SwinFusion seems to have richer details and brightness information. However, it does not retain some details from the infrared images, since this method utilizes a pixel maximum constraint loss for each pixel, see the red boxes in the fourth and fifth columns. In contrast, our ATFusion achieves a better trade-off between detail preservation and salient information retention.
@@
Table 2: Quantitative assessment of the compared approaches on the MSRS dataset (361 image pairs). The best and the second best performance is marked by bold and underlined, respectively.
Methods AG EN FMI $Q_{abf}$ SD SF SSIM VIFF
DenseFuse 2.0441 5.9331 0.9232 0.3651 7.4304 0.0235 0.9300 0.7053
U2Fusion 2.0000 5.8656 0.9134 0.3161 7.4645 0.0226 0.9120 0.6487
RFN-Nest 2.5485 6.2196 0.9229 0.5042 7.7855 0.0296 0.9554 0.7828
SwinFuse 1.9669 4.9449 0.8527 0.2038 5.7161 0.0262 0.9052 0.5980
SwinFusion 3.5414 6.6179 0.9312 0.6198 8.3973 0.0434 0.9695 1.0039
AEFusion 2.7009 6.5641 0.9088 0.4319 8.2501 0.0305 0.9293 0.8662
DATFuse 3.5227 6.5817 0.9085 0.5709 8.7157 0.0413 0.9394 0.7980
LRRNet 2.6309 6.3140 0.9097 0.3539 8.2313 0.0293 0.8872 0.7397
Ours 3.7619 6.5784 0.9318 0.7022 8.2459 0.0453 0.9759 1.0370
Tab. 2 displays the quantitative assessment of the compared approaches on 361 image pairs from the MSRS dataset. Our ATFusion method performs best in most metrics, except for EN and SD. In the EN case, it ranks third, scoring slightly less than the best. The reason for the relatively low value of the SD metric is that the DIIM module is used in our method, which makes the pixel distribution of the fused image relatively peaked. Overall, based on both the subjective and objective analyses, we can conclude that our ATFusion method achieves the best fusion performance and outperforms the competitors.
2) Results on the RoadScene Dataset: Fig. 6 shows five pairs of source infrared and visible images from the RoadScene dataset along with the fusion results obtained through the compared methods. Two localized areas in each image are magnified for better comparison. It is clear that the state-of-the-art methods used for comparison have some unsatisfactory performance. Indeed, results generated by DATFuse, AEFusion, U2Fusion, Densefuse, and SwinFusion have deficiencies when dealing with specific local information, see the red box in the first column and the green box in the second column. Besides, RFN-Nest fails to preserve well significant information from the infrared image in the fusion results, see the red and green boxes in the fifth columns. Although the fusion results of LRRNet and SwinFuse can highlight well the target information in the source image, they fail to transfer the detail information to the fused images in some scenes, see the red box in the third and fourth columns. By contrast, it can be seen that our ATFusion maintains the salient information in all results. This is because the proposed ATFusion employs the DIIM and ACIIM modules, which can generate results containing various types of information. Furthermore, the segmented pixel loss function enables our method to better balance the retention of detailed and salient information.
Figure 6: Qualitative assessment of the compared approaches on the RoadScene dataset.
Fig. 7 further reports quantitative results on the 20 image pairs of the RoadScene dataset by using the eight quality metrics. These results indicate that our method gets the best average values on four evaluation metrics (i.e., AG, SF, $Q_{abf}$, and FMI). In the case of the EN, VIFF, and SSIM quality metrics, it ranks second. Overall, our method achieves satisfactory objective performance, which demonstrates the crucial role of the proposed framework and the segmented pixel loss function.
Figure 7: Quantitative assessment of the compared approaches on the RoadScene dataset (20 image pairs). The average for all metrics is shown in the legend.
@@
3) Results on the RGB-NIR Scene Dataset: To further validate the effectiveness and generalization capabilities of our method, we directly utilize the trained model to test the RGB-NIR Scene dataset. Fig. 8 shows five image pairs from the RGB-NIR Scene dataset and the corresponding fusion results obtained by nine methods. We have magnified two local areas in each image for better visual comparison. It can be seen that the compared state-of-the-art methods still have some drawbacks compared to our approach. Specifically, SwinFusion, DATFuse, AEFusion, RFN-Nest, U2Fusion, and LRRNet fail to extract enough salient information from the infrared images compared to the other methods, see the green and red boxes in the second column. Moreover, the results produced by DenseFuse and SwinFuse are blurry in some local details, see the green box in the first and third columns. Compared with these methods, ATFusion cannot only preserve more detail, but also show a higher degree of target salience.
Figure 8: Qualitative assessment of the compared approaches on the RGB-NIR Scene dataset.
Figure 9: Quantitative assessment of the compared approaches on the RGB-NIR Scene dataset (20 image pairs). The average for all metrics is shown in the legend.
@@
Fig. 9 reports that our method achieves the best fusion performance for the EN, VIFF, $Q_{abf}$, and FMI metrics on the RGB-NIR Scene dataset. Instead, it is the second-best for the AG, SSIM, and SF metrics, with just a small gap compared to the best performance. A possible reason for the low SD value is that the pixels of the generated image are concentrated around the mean by using the DIIM module. In general, the ATFusion method achieves the best objective and subjective performance among the compared approaches on the 20 image pairs from the RGB-NIR Scene dataset, corroborating its excellent generalization ability.
4.4 Ablation Studies
@@
In this section, we have conducted two ablation studies, i.e., considering the network structure and the segmented pixel loss function.
@@
Figure 10: Fusion results obtained by three different network structures. (a) Infrared image, (b) visible image, (c) the results w/o DIIM, (d) the results w/o ACIIM, (e) Ours.
@@
1) Ablation Study on the Network Structure: To investigate the effectiveness of DIIM and ACIIM, we separately remove the DIIM (named “w/o DIIM"), and the ACIIM (named “w/o ACIIM"). We compare the fusion results obtained by w/o ACIIM, w/o DIIM, and the proposed method, as shown in Fig. 10. It can be seen that the fusion results produced by using w/o DIIM only contain detail information from the visible image, while the salient information from the infrared image is almost entirely lost, see the green and red boxes in Fig. 10(c). Besides, our fusion results have more detail information than the results by using the w/o ACIIM network, see the green and red boxes in Figs. 10(d)-(e). The main reason is that the proposed ATFusion embeds the DIIM and ACIIM modules, which have a good ability to capture the discrepancy information and the common information of the source images.
@@
Table 3: Quantitative assessment of the three different network structures. The best and the second best performance is marked by bold and underlined, respectively.
Methods AG EN FMI $Q_{abf}$ SD SF SSIM VIFF
w/o DIIM 3.4038 6.3734 0.9334 0.6028 8.1697 0.0419 0.9107 1.0234
w/o ACIIM 3.4836 6.5656 0.9310 0.7067 8.2095 0.0411 0.9791 0.9933
Ours 3.7619 6.5784 0.9318 0.7022 8.2459 0.0453 0.9759 1.0370
@@
Tab. 3 provides an objective comparison of our ATFusion with other two network structures. Our method can achieve the best performance for AG, EN, SD, SF, and VIFF, and the second-best performance for $Q_{abf}$, SSIM, and FMI. Therefore, the proposed DIIM and ACIIM modules can fully extract useful features, which enhance the retention of discrepancy information and the injection of common information in the fusion results.
@@
2) Ablation Study on the Pixel Loss Function: To identify the impact of the segmented pixel loss function, we conduct four experiments by using different pixel loss functions: $\mathcal{L}_{1}$ ($\alpha=100$), $\mathcal{L}_{2}$ ($\alpha=0$), $\mathcal{L}_{3}$ ($\alpha=50$), and $\mathcal{L}_{4}$ ($\alpha=80$) in Eq. (20). $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ correspond to the pixel maximum and average constraints, respectively. Fig. 11 shows the fusion outcomes with the five different loss functions. From the green boxes in Figs. 11(c), (e), and (f), it can be seen that the results generated by using $\mathcal{L}_{1}$, $\mathcal{L}_{3}$, and $\mathcal{L}_{4}$ have different degrees of detail loss. Besides, as the value of $\alpha$ increases, the degree of detail loss becomes more relevant. This is because the increasing value of $\alpha$ leads to more pixels in the image to be constrained by the maximum value selection loss, which brings to some information loss from one of the source images in fusion results. The results obtained by the $\mathcal{L}_{2}$ loss lack brightness information with respect to the results of the other loss functions, see the red box in Fig. 11(d). This is because the core of the $\mathcal{L}_{2}$ loss is the pixel average constraint. Anyway, setting $\alpha$ to 20 results in a satisfactory trade-off between maintaining detail and preserving saliency.
@@
Table 4: Quantitative assessment of the fusion results obtained by the proposed method with five different pixel loss functions. The best and the second best performance is marked by bold and underlined, respectively.
Loss Functions AG EN FMI $Q_{abf}$ SD SF SSIM VIFF
$\mathcal{L}_{1}$ ($\alpha=100$) 3.6773 6.5690 0.9316 0.7023 8.2338 0.0443 0.9729 1.0166
$\mathcal{L}_{2}$ ($\alpha=0$) 3.7080 6.5533 0.9315 0.7009 8.2287 0.0442 0.9774 1.0418
$\mathcal{L}_{3}$ ($\alpha=50$) 3.6805 6.5671 0.9315 0.7034 8.2197 0.0441 0.9750 1.0169
$\mathcal{L}_{4}$ ($\alpha=80$) 3.6027 6.5473 0.9308 0.7042 8.1918 0.0432 0.9731 1.0131
Ours ($\alpha=20$) 3.7619 6.5784 0.9318 0.7022 8.2459 0.0453 0.9759 1.0370
@@
Tab. 4 reports the quantitative assessment of our ATFusion exploiting various loss functions. The average outcomes for the several evaluation metrics are tabulated in Tab. 4. The best and second-best results for each metric are in boldface and underlined, respectively. It can be seen that the results generated by our method outperform the ones obtained by using the other loss functions. Therefore, setting $\alpha$ to 20 is a reasonable choice in our study.
4.5 Analysis of Some Configurations
@@
Figure 12: Fusion results obtained by the proposed method with different $\gamma$ values. (a) Infrared image, (b) visible image, (c)-(e) the qualitative results of different $\gamma$.
@@
Table 5: Quantitative assessment of the fusion results obtained by different $\gamma$ values. The best and the second best performance is marked by bold and underlined, respectively.
Hyperparameter AG EN FMI $Q_{abf}$ SD SF SSIM VIFF
$\gamma=0.5$ 3.6979 6.6916 0.9316 0.6918 8.3978 0.0440 0.9772 1.0281
$\gamma=0.75$ 3.7033 6.5727 0.9132 0.7007 8.2532 0.0443 0.9757 1.0213
$\gamma=1.0$ (Ours) 3.7619 6.5784 0.9318 0.7022 8.2459 0.0453 0.9759 1.0370
@@
1) Hyperparameter $\gamma$: As mentioned in Section III.C, the hyperparameter $\gamma$ controls the weight of the pixel and texture losses in the loss function. Based on our experience, we set $\gamma$ to 0.5, 0.75, and 1.0. Fig. 12 displays the fusion images produced by ATFusion using different $\gamma$ values in the loss function. It can be seen that the output obtained with $\gamma=1.0$ exhibits clear texture details, see the green box in Figs. 12(c)-(e). Tab. 5 presents an objective comparison of ATFusion using different $\gamma$ values. It is evident from the table that setting $\gamma$ to 1.0 yields the best performance.
@@
Table 6: Quantitative assessment of fusion results obtained by the proposed method with different numbers of feature fusion modules. The best and the second best performance is marked by bold and underlined, respectively.
Number AG EN FMI $Q_{abf}$ SD SF SSIM VIFF
Three 3.7129 6.4676 0.9316 0.6913 8.1105 0.0452 0.9707 1.0159
Two 3.6630 6.4268 0.9304 0.6900 8.0113 0.0444 0.9710 0.9844
One (Ours) 3.7619 6.5784 0.9318 0.7022 8.2459 0.0453 0.9759 1.0370
@@
2) Numbers of Feature Fusion Modules: We assembled one DIIM and two ACIIMs into a single feature fusion module, as shown in Fig. 1. The number of feature fusion modules can be decided by inspecting the fusion performance through objective metrics. Fig. 13 displays the fusion images produced by ATFusion using various numbers of feature fusion modules. It can be seen that using one feature fusion module in the proposed framework achieves the best overall effect, see the green and red boxes in Figs. 13(c)-(e). Besides, Tab. 6 reports the quantitative fusion results by using various numbers of feature fusion modules. We get the best results using one fusion model.
Figure 13: Fusion results obtained by the proposed method with five different pixel loss functions. (a) Infrared image, (b) visible image, (c) the results of two feature fusion modules, (d) the results of three feature fusion modules, (e) Ours.
4.6 Computational Efficiency